Cody Lind’s most recent credit card statement follows. His finance charge is 18% APR. Calculate Cody’s average daily balance, finance charge and new balance. (Round your final answers to the nearest cent.)

30-day billing cycle
9/2 Billing date $1,130 previous balance
9/7 Payment $380 cr.
9/13 Charge: Kohl’s $390
9/17 Payment $340 cr.
9/28 Charge: WalMart 125

Find...
1. Average daily balance: $

2. Finance charge: $

3. New balance: $
Answer #1

1. Average daily balance = $907

2. Finance charge = $13.61

3. New Balance = $938.61

Average Daily Balance:

$1130 x 5 Days = $5650

($1130 - 380) = 750 x 6 days = 4500

($750 + 390) = 1140 x 4 days = 4560

(1140 - 340) = $800 x 11 days = 8800

($800 + 125 ) = $925 x 4 days = 3700

$5650 + $4500 + $4560 + $8800 + $3700 = 27210

27210 / 30 days = $907

Average Daily Balance: $907

Finance Charge:

$27210 x .18 = $4898

30 x 12 = 360

4898 / 360 = 13.61

Finance Charge = 13.61

New Balance:

$1,130 - 380 + 390 - 340 + 125= $925

$925 + 13.61 = $938.61

New Balance = $938.61